# On a class of nonlinear BGK-type kinetic equations with density dependent collision rates

### v2

@@ Acknowledgements
J. Evans is supported by a Royal Society University Research Fellowship R1_251808 and before this a Leverhulme early career fellowship ECF-2021-134. For the purpose of open access, the author has applied a Creative Commons Attribution (CC-BY) licence to any Author Accepted Manuscript version arising from this submission. H. Yoldaş is supported by the Dutch Research Council (NWO) under the NWO-Talent Programme Veni ENW project MetaMathBio with the project number VI.Veni.222.288. Some part of this research was conducted while H. Yoldaş was visiting the Okinawa Institute of Science and Technology (OIST) through the Theoretical Sciences Visiting Program (TSVP).

### v3

@@ Acknowledgements
J. Evans is supported by a Royal Society University Research Fellowship R1_251808 and before this a Leverhulme early career fellowship ECF-2021-134. H. Yoldaş is supported by the Dutch Research Council (NWO) under the NWO-Talent Programme Veni ENW project MetaMathBio with the project number VI.Veni.222.288. Some part of this research was conducted while H. Yoldaş was visiting the Okinawa Institute of Science and Technology (OIST) through the Theoretical Sciences Visiting Program (TSVP). For the purpose of open access, the authors have applied a Creative Commons Attribution (CC-BY) licence to any Author Accepted Manuscript version arising from this submission.